Determine AGº for the phase change CO2(g) → CO2(aq) at 25°C. AH° = -19.4 kJ/mol AS° = +92.3 J/mol:K
Answer #1

ΔHo = -19.4 KJ/mol

ΔSo = 92.3 J/mol.K

= 0.0923 KJ/mol.K

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

ΔGo = ΔHo - T*ΔSo

ΔGo = -19.4 - 298.0 * 0.0923

ΔGo = -46.9 KJ/mol

Answer: -46.9 KJ/mol